OpenStudy (anonymous):

diff eqns: y=sinxcosx-cosx; y(0)=-1; show y is a soln of y'+tan(x)y'=cos^2(x). where y'=-sin^2(x)+sinx+cos^2(x).

OpenStudy (turingtest):

y'+tan(x)y=cos^2(x) ^ is that y supposed to have a ' on it or was that a typo?

OpenStudy (anonymous):

yes, both y in the diff eq should have a prime on it.

OpenStudy (turingtest):

well either way you just plug in the given solution and check that it is true: y'+tan(x)y'=cos^2(x) we are given that y'=-sin^2(x)+sinx+cos^2(x) now just sub that into the DE: -sin^2(x)+sinx+cos^2(x)+tan(x)(-sin^2(x)+sinx+cos^2(x))=cos^2(x) and check to see if it turns out to be true

OpenStudy (anonymous):

yes, i did but couldn't get it to work out. i used various trig ids to try to make it work....but could n't get it

OpenStudy (turingtest):

ok let me try...

OpenStudy (anonymous):

thank you!!

OpenStudy (turingtest):

well...\[-\sin^2x+\sin x+\cos^2x+\tan x(-\sin^2x+\sin x+\cos^2x)\neq\cos^2x\]according to the wolf

OpenStudy (anonymous):

huh. ok. well, that makes sense then that i couldnt' get it to work. thank you!!

OpenStudy (turingtest):

welcome, good luck finding your typo or whatever :P